OpenStudy (anonymous):

help with demoivre theorem (-1-i)^7

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

how do you solve this?

OpenStudy (misty1212):

step one is to write \[-1-i\] in trig form, polar form, or whatever you are calling it

OpenStudy (anonymous):

doesnt it get squared?

OpenStudy (misty1212):

no

OpenStudy (misty1212):

you have to raise it to the power of 7

OpenStudy (anonymous):

but with r=A^2+B^2

OpenStudy (anonymous):

help

OpenStudy (misty1212):

do you know how to write \[-1-i\] as \[r\left(\cos(\theta)+i\sin(\theta)\right)\]?

OpenStudy (misty1212):

that is what you have to do here

OpenStudy (misty1212):

if you cannot do that, i can help, but you can't do the problem without it

OpenStudy (amorfide):

laettner yes you do square A and B then square root it to get r

OpenStudy (misty1212):

\[r=\sqrt{a^2+b^2}\] yes which in this case is \(\sqrt2\) then you have to find \(\theta\)

OpenStudy (misty1212):

you know how to do that?